OpenStudy (anonymous):

0.5b+4=2(b+2) helppp

OpenStudy (anonymous):

The first thing you want to do is distribute the 2 into the (b+2) \[0.5b+4=2(b+2)\] \[0.5b+4=2(b)+2(2)\] Can you solve it from there?

OpenStudy (anonymous):

Um do i just subtract 2b-0.5b?like after

OpenStudy (anonymous):

yes, you will just move the b's to one side, and the numbers to the other. so you have: 2b-0.5b = 4-4 which in this case, b is just 0, but the next step is to divide each side by 1.5 to get b by itself

OpenStudy (anonymous):

oh oki.thankss